OpenStudy (anonymous):

Sketch the region enclosed by the given curves. y = 5 cos 4x, y = 5 − 5 cos 4x, 0 ≤ x ≤ π/4 Find it's area.

OpenStudy (anonymous):

This is a graph of the region

OpenStudy (anonymous):

I figured you had to go ahead and add the following integrations \[(\int\limits_{0}^{\Pi/12} (5\cos(4x)) - (5-5\cos(4x))dx ) + (\int\limits_{\Pi/12}^{\Pi/4} (5-5\cos(4x))-(5\cos(4x))dx)\] but every time I try it, the answer I get is wrong. Help !

OpenStudy (luigi0210):

Hello @beeca_boo73 Welcome to Openstudy :)

OpenStudy (anonymous):

Hi @Luigi0210 ! Thanks, I'm excited to discover this nifty site. Hopefully it'll be helpful. (:

OpenStudy (luigi0210):

@Psymon will assist you with your question.

OpenStudy (psymon):

Im already in 2 questions @Luigi0210

OpenStudy (luigi0210):

Btw if you have any questions or concerns about the site feel free to PM me, or anybody else with an A next to their name, hope you enjoy the site :)

OpenStudy (luigi0210):

Okay, I'll try and help.

OpenStudy (anonymous):

Okay, thanks. (: I'm just starting calc II and haven't had calc AB for a year and a half so I really can't remember how to do the integrations of trig functions in general.

OpenStudy (luigi0210):

Well okay, so you figured out where they intersect right?

OpenStudy (anonymous):

Yeah, I did. They intersect at \[\Pi/12\] I think I have the integration formula right I just don't know how to take the antiderivative of the cos functions.

OpenStudy (luigi0210):

\[\int\limits cosx dx=-sinx +C\] If I remember correctly

OpenStudy (luigi0210):

@Loser66 We have a student here waiting to learn :3

OpenStudy (luigi0210):

calc 2

OpenStudy (psymon):

Yeah, this is calc 2 stuff. Still in one problem, haha

OpenStudy (anonymous):

Yeah, calc II, semi rehash from calc AB.

OpenStudy (luigi0210):

@Psymon I thought you were busy?

OpenStudy (psymon):

So yeah, you need to determine a few things. the correct limits, the point of intersection and which function is higher and which function is lower. This will require two integrals because of us crossing and changing which function is higher. So before they intersect, which function is the higher of the two functions?

OpenStudy (psymon):

Nah, end the first integral at the intersection.

OpenStudy (luigi0210):

I think I see where you're going with this

OpenStudy (anonymous):

Before they intersect, the 5cos(4x) is on top, correct ? So the integration would be 5cos(4x) - (5-5cos(4x))

OpenStudy (psymon):

Correct.

OpenStudy (anonymous):

But then you have to swap them after that because 5cos(4x) is no longer on top.

OpenStudy (psymon):

Right. In which case you would do 5-5cos4x - 5cos4x for the second integral.

OpenStudy (anonymous):

And you integrate from the intersection to pi/4

OpenStudy (psymon):

The first integral you integrate up until the intersection. Then you add the second integral with limits starting at the intersection and ending at pi/4

OpenStudy (psymon):

\[\int\limits_{0}^{intersection}5\cos4x - (5-5\cos4x)dx + \int\limits_{intersection}^{\frac{ \pi }{ 4 }}5-5\cos4x-(5\cos4x)dx\]

OpenStudy (anonymous):

Right, I got that. I originally got pi/12 as the place they intersect. Was that correct or no?

OpenStudy (psymon):

Ill check.

OpenStudy (anonymous):

Thanks for trying @Loser66 ! (:

OpenStudy (psymon):

Yes, pi/12 is correct. : )

OpenStudy (psymon):

Therefore we have this :P \[\int\limits_{0}^{\frac{ \pi }{ 12 }}5\cos4x- (5-5\cos4x)dx + \int\limits_{\frac{ \pi }{ 12 }}^{\frac{ \pi }{ 4 }}5-5\cos4x-(5\cos4x)dx\]

OpenStudy (anonymous):

Yeah. Okay, great, so my set up is right. Yay! Now my main problem is I suck at integrating trig functions. So, I began by simplifying the integration to \[\int\limits_{0}^{\Pi/12}(10\cos(4x) - 5)dx + \int\limits_{\Pi/12}^{\Pi/4} (5-10\cos(4x))dx\] Does that look right?

OpenStudy (psymon):

Yes it does.

OpenStudy (psymon):

Now let's start with the first integral. Because cos is only to the first power and we have no wandering x's, we can simply say this: let u = 4x. therefore du = 4dx and dx = du/4. \[\int\limits_{}^{}10\cos(u)*\frac{ du }{ 4 }\]

OpenStudy (anonymous):

Okay. And can you pull the 10 to the front of the integral for that ? So when you integrate for cosu you'd get.. sinu/4 ?

OpenStudy (psymon):

Well you can factor out 10/4, but don't forget to put it back. So yes, cosine integrates to sin u where u = 4x. So what we end up with is: \[\frac{ 10\sin(u) }{ 4 }=\frac{ 5\sin(4x) }{ 2 } \]

OpenStudy (anonymous):

Okay, so you'd have 5sin(4x)/4 minus \[\int\limits_{}^{}5dx\] which would be 5x so for the first integration you'd end up with 5sin(4x)/2 - 5x from 0 - pi/12

OpenStudy (psymon):

We're working on that @Loser66 xD and yes, exactly. So we cansolve that and then get to our 2nd integral :3

OpenStudy (psymon):

?

OpenStudy (psymon):

We have two "intervals" though.

OpenStudy (anonymous):

sin(pi/3) is the square root of 2 over 2, right ?

OpenStudy (psymon):

sqrt(3)/2

OpenStudy (anonymous):

Aw, crap, okay, lol

OpenStudy (psymon):

Lol, we'll see if we run into any issues @Loser66 Thank you for watching over our work, though xD

OpenStudy (anonymous):

So the first integration would end up being \[5\sqrt{3}/4 - 5\sqrt{3}/12\] ?

OpenStudy (anonymous):

no wait minus 5(pi)/12

OpenStudy (psymon):

Lol, you caught yourself. Correct. Now we can work on adding the result of the 2nd interval : )

OpenStudy (anonymous):

Okay. 5 - 10cos(4x) so you'd separate it to the integral of 5dx, which would be 5x, minus the integral of 10cos(4x), which would end up being the same 5sin(4x)/2 right ?

OpenStudy (psymon):

Right, so we have 5x - 5sin(4x)/2 from pi/12 - pi/4

OpenStudy (anonymous):

Okay and you input the pi/4 for all the x's first, get that value and subtract the value from the expression when pi/12 is input for the x's after, right ?

OpenStudy (psymon):

Yep. Same thing we did from the first integral f(pi/4) - f(pi/12)

OpenStudy (anonymous):

so it would be (-5(pi)/4) - (5(sqrt 3)/4 - 5(pi)/12)

OpenStudy (anonymous):

?

OpenStudy (psymon):

5p/4 is positive.

OpenStudy (anonymous):

but I thought you'd subtract it from the 5sin(pi)/2 which would be zero ?

OpenStudy (anonymous):

Oh I put it in backwards

OpenStudy (psymon):

Lol, no worries.

OpenStudy (psymon):

We should have this: \[\frac{ 5\pi }{ 4 }-(\frac{ 5\pi }{ 12 }-\frac{ 5\sqrt{3} }{ 4 })\]

OpenStudy (anonymous):

I was using the formula from the first integration, my bad. okay so positive 5(pi)/4) so we'd get 5(pi)/6 - 5(sqrt 3)/4

OpenStudy (anonymous):

And add that to the first integrations answer of 5(sqrt 3)/4 - 5(pi)/12

OpenStudy (psymon):

Yes.

OpenStudy (psymon):

\[\frac{ 5\sqrt{3} }{ 4 }-\frac{ 5\pi }{ 12 }+\frac{ 5\pi }{ 6 }+\frac{ 5\sqrt{3} }{ 4 } \]

OpenStudy (anonymous):

so we'd get 5(sqrt3)/2 + 5(pi)/12 as the final answer ?

OpenStudy (psymon):

Yep xD

OpenStudy (anonymous):

THANK THE LORD You are a beautiful person, thanks so much. (:

OpenStudy (psymon):

Yep np ^_^